Use the following information for questions 1-7. A corporation has 9,000,000 shares of stock outstanding at a price of $40 per share. Net Income is $28,000,000 and they just paid a dividend of $2 and the dividend is expected to grow by 5% per year forever (Therefore next year’s dividend will be 2*(1.05) = $2.10). The stock has a beta of .9, the current risk free rate is 4%, and the market risk premium is 6%. The corporation also has 300,000 bonds outstanding with a price of $1,100 per bond. The bond has a coupon rate of 8% with semiannual interest payments, a face value of $1,000, and 13 years to go until maturity. The company plans on adding debt until they reach their target debt ratio of 70%. They expect their before-tax cost of debt to be 9% and their cost of equity to be 14% under this new capital structure. The tax rate is 25%

5. What is their WACC using their target capital structure and expected costs of debt and equity?

a) 7.7%          b) 8.9%            c) 9.4%            d) 10.2%

6. Given the new cost of debt, what should be the new price of the bond?

a) $925          b) $960            c) $1,025         d) $1,175

7. Given the new cost of equity, what should be the new price of the stock?

a) $23.33       b) $27.25         c) $33.50         d) $36.67
Answer #1

5) Cost of debt * Weight of debt* (1- tax rate) + Cost of equity * Weight of equity WACC = Cost of debt 9% weight of debt 0.7

Face value of the bond $1000 6) Coupon 1000*0.08/2 = 40 Number of payments = 13*2 = 26 Interest rate = 9%/2 = 4.5% %3D Price

7) The new price of stock next year dividend/( cost of equity - growth rate) Next year dividend Cost of equity 2.1 14% Growth